# Tiled Flash Linear Attention: More Efficient Linear RNN and xLSTM Kernels

### v2

@@ F.2 Simplified FLOP Count Summary for the mLSTM
Tables 14, 15 and 16 summarize the simplified FLOP counts for the chunkwise-parallel, fully parallel and recurrent mLSTM formulation.
@@
FLOPs mLSTMexp mLSTMsig
Recurrent computation of the inter chunk states
Gates: $50.5L^{2}+6.5L+$ 5 $10.5L^{2}+7.5L+$ 1
Numerator: $2d_{qk}d_{hv}+2Ld_{qk}d_{hv}+Ld_{qk}$ $2d_{qk}d_{hv}+2Ld_{qk}d_{hv}+Ld_{qk}$
Denominator: $2d_{qk}+2Ld_{qk}$
Parallel computation of the intra chunk outputs
Cumulative Forget Gates: $0.5L^{2}+2.5L$ $0.5L^{2}+2.5L$
Gate Matrix: $F_{\text{causal}}\times\left(5L^{2}+2L\right)$ $F_{\text{causal}}\times 3L^{2}$
Intra Outputs: $F_{\text{causal}}\times\left(2L^{2}(d_{qk}+d_{hv})+3L^{2}\right)$ $F_{\text{causal}}\times\left(2L^{2}(d_{qk}+d_{hv})+3L^{2}\right)$
Parallel computation of the inter chunk outputs
Inter Outputs: $2Ld_{qk}d_{hv}+3Ld_{qk}$ $2Ld_{qk}d_{hv}+Ld_{qk}$
Combination of inter and intra chunk outputs
Output Combination: $2Ld_{hv}+4L$ $Ld_{hv}$
@@
FLOPs mLSTMexp mLSTMsig Difference chunkwise-parallel: $TLF_{\text{causal}}\left(2(d_{qk}+d_{hv})+8\right)+TL+2TF_{\text{causal}}$ $+T\left(4d_{qk}d_{hv}+6d_{qk}+4d_{hv}+13\right)$ $+\frac{T}{L}\left(2d_{qk}d_{hv}+2d_{qk}+5\right)$ $TLF_{\text{causal}}\left(2(d_{qk}+d_{hv})+6\right)+TL$ $+T\left(4d_{qk}d_{hv}+2d_{qk}+d_{hv}+11\right)$ $+\frac{T}{L}\left(2d_{qk}d_{hv}+5\right)$ $2TLF_{\text{causal}}+2TF_{\text{causal}}$ $+T\left(4d_{qk}+3d_{hv}+2\right)$ $+2\frac{T}{L}d_{qk}$ fully parallel: $T^{2}F_{\text{causal}}\left(2(d_{qk}+d_{hv})+6\right)$ $+2TF_{\text{causal}}+6.5T^{2}+2.5T$ $T^{2}F_{\text{causal}}\left(2(d_{qk}+d_{hv})+2\right)$ $+6.5T^{2}+4.5T$ $4T^{2}F_{\text{causal}}$ $+2TF_{\text{causal}}-2.5T$ recurrent: $T\left(6d_{qk}d_{hv}+7d_{qk}+d_{hv}+12\right)$ $T\left(6d_{qk}d_{hv}+d_{qk}+2\right)$ $T\left(6d_{qk}+d_{hv}+10\right)$
@@ F.4 FLOP Comparison between different mLSTM Formulations
Figure 18: FLOP counts for the recurrent, fully parallel and chunkwise-parallel mLSTM formulation of mLSTMsig. Left: $d_{qk}$=256, $d_{hv}$=512. Middle: $d_{qk}$=128, $d_{hv}$=256 Right: $d_{qk}$=64, $d_{hv}$=128. We count the number of FLOPs for a one head and one sequence of length $T$=8192 for different head dimensions and vary the chunk size $L$. By varying the chunk size $L$ the chunkwise-parallel formulation FLOP counts transition between the recurrent and fully parallel FLOP counts. Smaller head dimensions decrease the overall FLOP count.
@@ F.5 FLOP-Optimal Chunk Size
In order to minimize the FLOPs to compute, $L_{\text{opt,FLOP}}$ indicates we should use rather small chunk sizes (e.g. $L$=16 for $d_{hv}$=512). However, the chunk size $L$ does not only affect the FLOPs but also the memory IO (for e.g. for loading and storing the memory cell states). Therefore, in order to find the optimal chunksize that minimizes the runtime, we need to consider the memory IO as well, which we do in the next section (see Appendix G).
@@ G.1 Theoretical Runtime
where we assume that the queries, keys, values and input and forget gate are stored in bfloat16, i.e. $\text{bytes}_{if}$=$\text{bytes}_{qkv}$=2.
@@
We show the theoretical runtime of our mLSTMsig kernel for 7B model size (i.e. $d_{hv}$=512, $p_{qk}$=0.5, $N_{\text{head}}$=8) for $N_{\text{batch}}$=8, $T$=8192 and $\text{bytes}_{Cmn}$=4 on different NVIDIA A100, H100 and B200 GPUs (according to Table 18) in Figure 20. We observe that newer GPUs (e.g. H100 or B200) are faster and have a higher memory bandwidth, which results in a lower runtime. Moreover, for newer GPUs, the runtime becomes less sensitive to the chunk size $L$, as the curve becomes flatter. Moreover, there exists an optimal chunk size that minimizes the runtime, which is determined by the physical constraints of the hardware. This optimal chunk size increases on more recent GPUs (e.g. B200). We explore the runtime-optimal chunk size more in depth in Appendix G.4.
Figure 20: Theoretical Runtime of TFLA mLSTMsig Forward Pass with varying Chunk Size $L$ on different accelerators. We plot the theoretical runtime of mLSTMsig with 7B size (i.e. $d_{hv}$=512, $p_{qk}$=0.5, $N_{\text{head}}$=8) for $N_{\text{batch}}$=8, $T$=8192 and $\text{bytes}_{Cmn}$=4 on NVIDIA A100, H100 and B200 GPUs (according to Tab. 18). We also plot the measured runtime of the mLSTMsig kernel on NVIDIA H100. Newer GPUs (e.g. H100, B200) are faster and have a higher memory bandwidth, which results in a lower theoretical runtime.
@@ G.2 Arithmetic Intensity
Then, (1) if  $\mathcal{I}_{\text{ algo}}>\mathcal{I}_{\text{ acc}}$, the kernel is likely to be compute-bound and (2) if  $\mathcal{I}_{\text{ algo}}<\mathcal{I}_{\text{ acc}}$, the kernel is likely memory-bound. The accelerator arithmetic intensity  accsubscript $\mathcal{I}_{\text{ acc}}$ is the minimum arithmetic intensity required to achieve maximum performance of the accelerator.
@@
We plot the arithmetic intensity of the TFLA mLSTMsig forward pass over the chunk size $L$ in Figure 21 and vary the head dimension $d_{hv}$ and the precision of the memory cell states $\text{bytes}_{Cmn}$. Additionally, we indicate the arithmetic intensity  accsubscript $\mathcal{I}_{\text{ acc}}$ for different NVIDIA GPUs (according to Table 18). Values above the accelerator arithmetic intensity indicate that the kernel is likely to be compute-bound for the corresponding chunk sizes, while values below indicate that the kernel is likely memory-bound.
@@ G.3 Roofline Analysis
If we then plot the arithmetic intensity of a kernel as a column that hits the roof, either it hits the flat part of the roof, meaning performnace is compute-bound or performance is ultimately memory bound (Williams et al., 2009).
@@ G.4 Runtime-Optimal Chunk Size
The main advantage of TFLA is that we can choose arbitrary chunk sizes $L$ to trade off between FLOPs and memory IO and minimize the runtime (see Figure 6 and 23). The reason for this tradeoff is that the chunk size $L$ affects the number of FLOPs and the memory IO (e.g. for loading and storing the memory cell states) (see Appendix F). In this section, we use this insight and our theoretical runtime model from Appendix G.1 to determine the theoretical runtime-optimal chunk size for the TFLA mLSTMsig forward kernel.
@@
where $\mathcal{I}_{\text{acc}}=\frac{\alpha_{\text{ FLOPs}{{}}}}{\beta_{\text{ Bytes}{{}}}}$ is the accelerator arithmetic intensity from Equation (119). Compared to the FLOP-optimal chunk size $L_{\text{opt,FLOP}}$ (see Equation (113)), the runtime-optimal chunk size $L_{\text{opt,Runtime}}$ additionally depends on the arithmetic intensity $\mathcal{I}_{\text{acc}}$ of our hardware and the precision of our states $\text{bytes}_{Cmn}$. The runtime-optimal chunk size grows proportional to the square root of the head dimension $d_{hv}$ (i.e. $\mathcal{O}(\sqrt{d_{hv}})$) and the accelerator arithmetic intensity $\mathcal{I}_{\text{acc}}$ (i.e. $\mathcal{O}(\sqrt{\mathcal{I}_{\text{acc}}})$). We visualize these trends in Figure 24 and 25.

### v3

@@ F.2 Simplified FLOP Count Summary for the mLSTM
Tables 15, 16 and 17 summarize the simplified FLOP counts for the chunkwise-parallel, fully parallel and recurrent mLSTM formulation.
@@
FLOPs mLSTMexp mLSTMsig Recurrent computation of the inter chunk states Gates: $0.5L^{2}+6.5L+5$ $0.5L^{2}+7.5L+1$ Numerator: $2d_{qk}d_{hv}+2Ld_{qk}d_{hv}+Ld_{qk}$ $2d_{qk}d_{hv}+2Ld_{qk}d_{hv}+Ld_{qk}$ Denominator: $2d_{qk}+2Ld_{qk}$ Parallel computation of the intra chunk outputs Cumulative Forget Gates: $0.5L^{2}+2.5L$ $0.5L^{2}+2.5L$ Gate Matrix: $F_{\text{causal}}\times\left(5L^{2}+2L\right)$ $F_{\text{causal}}\times 3L^{2}$ Intra Outputs: $F_{\text{causal}}\times\left(2L^{2}(d_{qk}+d_{hv})+3L^{2}\right)$ $F_{\text{causal}}\times\left(2L^{2}(d_{qk}+d_{hv})+3L^{2}\right)$ Parallel computation of the inter chunk outputs Inter Outputs: $2Ld_{qk}d_{hv}+3Ld_{qk}$ $2Ld_{qk}d_{hv}+Ld_{qk}$ Combination of inter and intra chunk outputs Output Combination: $2Ld_{hv}+4L$ $Ld_{hv}$
@@
FLOPs mLSTMexp mLSTMsig Difference chunkwise-parallel: $TLF_{\text{causal}}\left(2(d_{qk}+d_{hv})+8\right)+TL+2TF_{\text{causal}}$ $T\left(4d_{qk}d_{hv}+6d_{qk}+4d_{hv}+13\right) +$$\frac{T}{L}\left(2d_{qk}d_{hv}+2d_{qk}+5\right)$ $TLF_{\text{causal}}\left(2(d_{qk}+d_{hv})+6\right)+TL$ $T\left(4d_{qk}d_{hv}+2d_{qk}+d_{hv}+11\right) +$$\frac{T}{L}\left(2d_{qk}d_{hv}+5\right)$ $2TLF_{\text{causal}}+2TF_{\text{causal}}$ $T\left(4d_{qk}+3d_{hv}+2\right) +$$2\frac{T}{L}d_{qk}$ fully parallel: $T^{2}F_{\text{causal}}\left(2(d_{qk}+d_{hv})+6\right)$ $2TF_{\text{causal}}+6.5T^{2}+2.5T$ $T^{2}F_{\text{causal}}\left(2(d_{qk}+d_{hv})+2\right)$ $6.5T^{2}+4.5T$ $4T^{2}F_{\text{causal}}$ $2TF_{\text{causal}}-2.5T$ recurrent: $T\left(6d_{qk}d_{hv}+7d_{qk}+d_{hv}+12\right)$ $T\left(6d_{qk}d_{hv}+d_{qk}+2\right)$ $T\left(6d_{qk}+d_{hv}+10\right)$
@@ F.4 FLOP Comparison between different mLSTM Formulations
Figure 19: FLOP counts for the recurrent, fully parallel and chunkwise-parallel mLSTM formulation of mLSTMsig. Left: $d_{qk}$=256, $d_{hv}$=512. Middle: $d_{qk}$=128, $d_{hv}$=256 Right: $d_{qk}$=64, $d_{hv}$=128. We count the number of FLOPs for a one head and one sequence of length TT=8192 for different head dimensions and vary the chunk size $L$. By varying the chunk size $L$ the chunkwise-parallel formulation FLOP counts transition between the recurrent and fully parallel FLOP counts. Smaller head dimensions decrease the overall FLOP count.
@@ F.5 FLOP-Optimal Chunk Size
In order to minimize the FLOPs to compute, $L_{\text{opt,FLOP}}$ indicates we should use rather small chunk sizes (e.g. LL=16 for $d_{hv}$=512). However, the chunk size $L$ does not only affect the FLOPs but also the memory IO (for e.g. for loading and storing the memory cell states). Therefore, in order to find the optimal chunksize that minimizes the runtime, we need to consider the memory IO as well, which we do in the next section (see Appendix G).
@@ G.1 Theoretical Runtime
where we assume that the queries, keys, values and input and forget gate are stored in bfloat16, i.e. $\text{bytes}_{if}$$\text{bytes}_{qkv}$=2.
@@
We show the theoretical runtime of our mLSTMsig kernel for 7B model size (i.e. $d_{hv}$=512, $p_{qk}$=0.5, $N_{\text{head}}$=8) for $N_{\text{batch}}$=8, TT=8192 and $\text{bytes}_{Cmn}$=4 on different NVIDIA A100, H100 and B200 GPUs (according to Table 19) in Figure 21. We observe that newer GPUs (e.g. H100 or B200) are faster and have a higher memory bandwidth, which results in a lower runtime. Moreover, for newer GPUs, the runtime becomes less sensitive to the chunk size $L$, as the curve becomes flatter. Moreover, there exists an optimal chunk size that minimizes the runtime, which is determined by the physical constraints of the hardware. This optimal chunk size increases on more recent GPUs (e.g. B200). We explore the runtime-optimal chunk size more in depth in Appendix G.4.
Figure 21: Theoretical Runtime of TFLA mLSTMsig Forward Pass with varying Chunk Size $L$ on different accelerators. We plot the theoretical runtime of mLSTMsig with 7B size (i.e. $d_{hv}$=512, $p_{qk}$=0.5, $N_{\text{head}}$=8) for $N_{\text{batch}}$=8, TT=8192 and $\text{bytes}_{Cmn}$=4 on NVIDIA A100, H100 and B200 GPUs (according to Tab. 19). We also plot the measured runtime of the mLSTMsig kernel on NVIDIA H100. Newer GPUs (e.g. H100, B200) are faster and have a higher memory bandwidth, which results in a lower theoretical runtime.
@@ G.2 Arithmetic Intensity
Then, (1) if  $\mathcal{I}_{\text{ algo}}>\mathcal{I}_{\text{ acc}}$, the kernel is likely to be compute-bound and (2) if  $\mathcal{I}_{\text{ algo}}<\mathcal{I}_{\text{ acc}}$, the kernel is likely memory-bound. The accelerator arithmetic intensity  $\mathcal{I}_{\text{ acc}}$ is the minimum arithmetic intensity required to achieve maximum performance of the accelerator.
@@
We plot the arithmetic intensity of the TFLA mLSTMsig forward pass over the chunk size $L$ in Figure 22 and vary the head dimension $d_{hv}$ and the precision of the memory cell states $\text{bytes}_{Cmn}$. Additionally, we indicate the arithmetic intensity  $\mathcal{I}_{\text{ acc}}$ for different NVIDIA GPUs (according to Table 19). Values above the accelerator arithmetic intensity indicate that the kernel is likely to be compute-bound for the corresponding chunk sizes, while values below indicate that the kernel is likely memory-bound.
@@ G.3 Roofline Analysis
If we then plot the arithmetic intensity of a kernel as a column that hits the roof, either it hits the flat part of the roof, meaning performnace is compute-bound or performance is ultimately memory bound Williams et al. (2009).
@@ G.4 Runtime-Optimal Chunk Size
The main advantage of TFLA is that we can choose arbitrary chunk sizes $L$ to trade off between FLOPs and memory IO and minimize the runtime (see Figure 6 and 24). The reason for this tradeoff is that the chunk size $L$ affects the number of FLOPs and the memory IO (e.g. for loading and storing the memory cell states) (see Appendix F). In this section, we use this insight and our theoretical runtime model from Appendix G.1 to determine the theoretical runtime-optimal chunk size for the TFLA mLSTMsig forward kernel.
@@
where $\mathcal{I}_{\text{acc}}=\frac{\alpha_{\text{ FLOPs}{{}}}}{\beta_{\text{ Bytes}{{}}}}$ is the accelerator arithmetic intensity from Equation (109). Compared to the FLOP-optimal chunk size $L_{\text{opt,FLOP}}$ (see Equation (103)), the runtime-optimal chunk size $L_{\text{opt,Runtime}}$ additionally depends on the arithmetic intensity $\mathcal{I}_{\text{acc}}$ of our hardware and the precision of our states $\text{bytes}_{Cmn}$. The runtime-optimal chunk size grows proportional to the square root of the head dimension $d_{hv}$ (i.e. $\mathcal{O}(\sqrt{d_{hv}})$) and the accelerator arithmetic intensity $\mathcal{I}_{\text{acc}}$ (i.e. $\mathcal{O}(\sqrt{\mathcal{I}_{\text{acc}}})$). We visualize these trends in Figure 25 and 26.